OpenStudy (anonymous):

how to integrate 1/(x^2+10*x+6)

OpenStudy (anonymous):

partial fractions is my bet.

OpenStudy (anonymous):

you want the answer or the method?

OpenStudy (anonymous):

just checked with wolfram and they give method as some sort of completing the square, getting arctanh and the converting back, but i would use partial fractions. set the denominator= and solve

OpenStudy (anonymous):

you get \[x^2+10x+6=0\] \[x=-5+\sqrt{19},x=-5-\sqrt{19}\] so you have \[\int\frac{dx}{(x-(-5+\sqrt{19})(x-(-5-\sqrt{19})}\]

OpenStudy (anonymous):

then solve \[\frac{A}{x-(-5+\sqrt{19})}+\frac{B}{x-(-5-\sqrt{19})}\] for A and B\]

OpenStudy (anonymous):

hmm maybe be easier way but i don't see it. not that bad actually but the calculation stinks

OpenStudy (anonymous):

oh no it doesn't it is easy. \[B=\frac{1}{2\sqrt{19}}\] and \[A=-\frac{1}{2\sqrt{19}}\]

OpenStudy (anonymous):

integrate \[-\frac{1}{2\sqrt{19}}\int\frac{dx}{x+5-\sqrt{19}}\] get \[-\frac{1}{2\sqrt{19}}\ln(x+5-\sqrt{19})\] and similarly \[\frac{1}{2\sqrt{19}}\int \frac{dx}{x+5+\sqrt{19}}=\frac{1}{2\sqrt{19}}\ln(x+5+\sqrt{19})\]